OpenStudy (anonymous):

SOMEONE PLEASE HELP ME WITH THIS ALGEBRA 2 . its hard and im really stressed

OpenStudy (jhannybean):

Well, whats the question? :D

OpenStudy (anonymous):

Match the complex expression to its correct simplified form. 1. i^22 2. i&7 3. 5 + 8i - 4 + 2i 4. (2 - 3i)(3 + 5i) 5. (2 - i)(2 + i)

OpenStudy (astrophysics):

i^22 = -1

OpenStudy (usukidoll):

hmm... well for the first one , we know that \[i^2= -1 \] so we can split the \[i^{22} \] with the exponent rule but that can take long... wow @Astrophysics so fast lol

OpenStudy (anonymous):

lol . i just dont understand it period

OpenStudy (usukidoll):

well I know why it was done fast.. the exponent was an even number XD it's like writing i^2 11 times

OpenStudy (astrophysics):

Main thing is you understand that i^2 = -1

OpenStudy (astrophysics):

Which can be proved from \[i^2 = (0+i)^2 \implies (0+i)(0+i) \implies ...-1+0i = -1\]

OpenStudy (usukidoll):

long way... not recommended lol \[(i^2)(i^2)(i^2)(i^2)(i^2)(i^2)(i^2)(i^2)(i^2)(i^2)(i^2)\] \[(-1)(-1)(-1)(-1)(-1)(-1)(-1)(-1)(-1)(-1)(-1)\] \[(1)(1)(1)(1)(1)(-1) \rightarrow (1)(-1)= -1 \]

OpenStudy (astrophysics):

lool

OpenStudy (astrophysics):

Nice haha

OpenStudy (usukidoll):

i^{100} is gonna hurt by hand XD

OpenStudy (usukidoll):

so next question For 3. 5 + 8i - 4 + 2i we can combine like terms so our answer is in the form of ai+b or a+bi

OpenStudy (astrophysics):

\[i^7 \implies (i^2)(i^2)(i^2)i\]

OpenStudy (anonymous):

so number 2 would be i^7 (-1)(-1)(-1)(-1)(-1)(-1)(-1) = -1

OpenStudy (astrophysics):

Naw, \[i = \sqrt{-1}\]

OpenStudy (usukidoll):

oh I read that second problem wrong. I thought it was i and 7 XD

OpenStudy (astrophysics):

\[i^7 = i^6 i = -1 \times i = -i\]

OpenStudy (anonymous):

okay so problem 3 would be i^3 ?

OpenStudy (usukidoll):

problem 3 has 3. 5 + 8i - 4 + 2i we can combine like terms so our answer is in the form of ai+b or a+bi

OpenStudy (astrophysics):

Yeah just combine like terms haha

OpenStudy (anonymous):

-1+10i

OpenStudy (astrophysics):

close but should be +1

OpenStudy (usukidoll):

we can rearrange it to make it a bit easier 5-4+8i+2i -> 1+10i

OpenStudy (usukidoll):

well the attempt wasn't that bad.... just 5-4=1

OpenStudy (usukidoll):

the last 2 problems is just using FOIL and using the fact that\[ i^2 = -1 \] again the final form of these questions is a+bi or ai+b

OpenStudy (anonymous):

(2 - i)(2 + i) 4 + 2i - 2i - i^2 4+i^2

OpenStudy (anonymous):

did i do that right

OpenStudy (astrophysics):

\[(2-i)(2+i) \implies (2)(2)+2i+(2)(-i)-i^2 \implies 4+2i-2i-i^2 = 4-i^2 = 4-(-1) = 4+1 = 5\]

OpenStudy (astrophysics):

\[4-i^2 = 4-(-1) \implies 4+1 = 5\]

OpenStudy (usukidoll):

(2-i)(2+i) 4+2i-2i-i^2 4-i^2 since i^2 = -1 4-(-1) = 4+1 = 5

OpenStudy (astrophysics):

Sorry it got cut off there

OpenStudy (astrophysics):

You're on the right path though, just a sign error :-)

OpenStudy (usukidoll):

it was a good attempt though .. just when we see i^2, we need to use substitution Replace i^2 with -1

OpenStudy (anonymous):

(2 - 3i)(3 + 5i) 6 + 10i - 9i - 15i 6 - 14i

OpenStudy (astrophysics):

Hmm I think not

OpenStudy (usukidoll):

(2-3i)(3+5i) 6+10i-9i-15i^2

OpenStudy (usukidoll):

remember... i^2 = -1

OpenStudy (anonymous):

how do you apply -15 and i^2

OpenStudy (astrophysics):

\[(2-3i)(3+5i) = 6+10i-9i-15i^2 \implies 6+i-15i^2\] \[6+i-15(-1) \implies 6+i+15 \implies 21 +i\]

OpenStudy (astrophysics):

Do you see your error now?

OpenStudy (usukidoll):

i^2 = -1 replace the i^2 with -1 6+10i-9i-15i^2 since i^2 = -1 6+10i-9i-15(-1)

OpenStudy (usukidoll):

6+i+15 21+i

OpenStudy (astrophysics):

\[\huge \checkmark\]

OpenStudy (anonymous):

omg thank yall . !!!!!! can you help me with something else ?

OpenStudy (anonymous):

1. X^2 + 9 = 0 2. 3x^2 + 11 = -181

OpenStudy (usukidoll):

are these questions solve for x?

OpenStudy (astrophysics):

You can set x^2 = -9 this has to do with imaginary numbers as well

OpenStudy (anonymous):

Solve the following: x^2 + 9 = 0 Select one: a. 2 + 4i b. + 3i c. not possible d. + 9i

OpenStudy (astrophysics):

Remember as I mentioned before \[i = \sqrt{-1}\]

OpenStudy (anonymous):

Solve: 3x^2 + 11 = -181 Select one: a. not possible b. + 8i c. 1 + 13i d. + 4i

OpenStudy (astrophysics):

\[x^2 = - 9\] what do you get

OpenStudy (usukidoll):

ah yes once we square root both sides on x^2=-9 we're gonna have an imaginary case. since negative signs in radicals aren't allowed, the result is imaginary.. denoted by i

OpenStudy (anonymous):

3

OpenStudy (anonymous):

no

OpenStudy (anonymous):

3i ?

OpenStudy (astrophysics):

That's one solution remember when we take a square root we get a positive and negative value :-)

OpenStudy (usukidoll):

hmm I guess for this multiple choice, it's only considering the positive version when taking the square root